The producer of a downloadable antivirus software program spends exactly ​$2,750,000 producing the first copy and incurring various costs required to make the software​ "user-friendly." The firm can produce and distribute additional copies at a​ per-unit cost of ​$2.00. If the company sold as many copies as consumers wished to purchase at a price of ​$2.00 per​ copy, it would sell 425,000 copies. If the company maximizes its economic profits in the​ short-run, it sells 225, 000 copies at a price of ​$45. ​Finally, the company earns zero economic profits when it sells 275,000 copies. What are the​ firm's economic profits​ (or losses) if it sells 425,000 copies of the antivirus software program at a ​$2.00 price per​ copy? ​$ . What are the maximum economic profits that the firm can earn in the short​ run? ​$ ▼ . What is marginal revenue when the firm maximizes its​ short-run economic​ profits? ​$ In the long​ run, after entry of competing​ firms, to the nearest​ dollar, and including the correct​ sign, what amount of economic profits will this firm​ earn? ​$
Answer #1

Per Unit marginal Cost= 2, Initial Cost = 2.75 million.

1. If it sells 425,000 copies: Price = MC= $2, Revenue= P.Q = 2*425,000 = Total Variable Cost for the firm.
Hence, at this level, it makes no profits and only manages to  stay afloat!
Total Loss= 2.75 million

2. Maximising Profit occurs at: 225,000 copies at $45 each.
Hence, at this point, MR= MC= 2
Hence, Total Revenue= 12.375 million. Total Cost = 2.75 million + 2*225,000 = 3.2 million
Maximum profits= 9.175 million

3. Zero Economic Profits occurs at 275,000 copies
Hence, 2.75 million + 275000*2 = P*275000
or P = 10+ 2 = $12

3. In the long run, firm makes only normal profits
Hence, the firm makes Zero Economic Profits (above case)

Let me know in case of any queries or if I have made any mistakes